OpenStudy (anonymous):

10=-2/3(4x+5) please solve and show the steps Thank You so much

OpenStudy (saifoo.khan):

do you want to solve that for x?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

30 = -2(4x+5) 30/2 = 4x + 5 4x + 5 = 15 4x = 15 - 5 4x = 10 i think u can solve now! ;)

OpenStudy (anonymous):

thats correct unless 4x+5 was supposed to be in the denominater

OpenStudy (anonymous):

thank you so much you guys

OpenStudy (saifoo.khan):

np.

OpenStudy (anonymous):

np